40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2512     Entry time: Wed Jan 13 12:01:06 2010
Author: Alberto 
Type: Update 
Category: Computers 
Subject: c1dcuepics, c1lsc rebooted this morning 
Since last night the alignemtn scripts couldn't work.
c1lsc wasn't working properly because attempts to lock the X arm would try to control ETMY and attempts of locking the Y arm, wouldn't actuate any optics.
Also, another sign of a malfunctioning c1lsc was that one of the LSC filter modules, FM6, couldn't get loaded properly. It looked like only half loaded on the LSC MEDM screen.
On the other hand, plotting the trend of the last month, c1lsc's CPU didn't look more loaded that usual.
 
Rebooting and restarting C1lsc wasn't enough and I also had to reboot c1dcuepics a couple of times beforse getting things back to work.
ELOG V3.1.3-